THE ROTATIONAL SPECTRUM OF CIS- AND TRANS-FORMIC ACID, REVISITED

\underline{BRENDA P. WINNEWISSER}^{a}, MANFRED WINNEWISSER,^{b} MONIKA STEIN, Physikalisch-Chemisches Institut der Justus-Liebig-Universität, Heinrich-Buff-Ring 58, D-35392 Gießen, Germany; MANFRED BIRK, GEORG WAGNER, DLR Oberpfaffenhofen, Postfach 11 16, D-82230 Wessling, Germany; GISBERT WINNEWISSER, KOICHI M. T. YAMADA, SERGEI P. BELOV, I. Physikalisches Institut, Universität zu Köln, Zülpicherstrasse 77, D-50937 Köln, Germany; OLEG I. BASKAKOV, Department of Quantum Radiophysics, Kharkov State University, Svobody Sq. 4, 310077 Kharkov, Ukraine.

The initial experimental confirmation of the existence of the cis rotamer of formic acid, HCOOH, was based on a small number of pure rotational transitions in the microwave region [W. H. Hocking, Z. Naturforsch. \underline{31a}, 1113–1121 (1976)]. Since then, there have been no further report of the observation of spectral features attributable to this species. In order to extend the study of the spectrum of cis-HCOOH we have recorded the far infrared spectrum of HCOOH from 30 to 100 cm^{−1} (Oberpfaffenhofen). This spectrum has now been supplemented by millimeter wave measurements (Kharkov) and submillimeter wave measurements (Köln), so that a very extensive set of data is now available for both rotamers. The analysis of the spectrum (Giessen, Kharkov)has now yielded a more complete set of ground state constants for each rotamer than was previously available, providing in particular a sound basis for the search for infrared transitions of cis-HCOOH.

aPresent address: Department of Physics, The Ohio State University, Columbus, Ohio 43210

bPresent address: Department of Physics, The Ohio State University, Columbus, Ohio 43210